# On the foundations and applications of Lorentz-Finsler Geometry

### v1

@@ 3.2.2. The case with timelike boundary
As seen in §3.1, relevant properties for globally hyperbolic spacetimes can be extended to the case of spacetimes-with-timelike-boundary and, then, to the Lorentz-Finsler setting. The recent reference [84] considers both, Lorentz-Finsler metrics and spacetimes with timelike boundary and, although the outcomes are equal in the Lorentz and the Lorentz-Finsler cases, the proof in the latter becomes subtler.
The first natural question on $\mathcal{N}$ is its Haudorffness, only ensured above by global hyperbolicity. The step in the causal ladder just below global hyperbolicity is causal simplicity, that is, being causal with all $J^{+}(p),J^{-}(p)$ closed. However, this is not enough to ensure Hausdorffnes, as a counterxample by Hedicke and Suhr [81] shows (see a simpler one in [84, §5.2] and further results in [80]). However, in the case of timelike boundary, both causal simplicity and Hausdorffness of $\mathcal{N}$, are ensured by the (infinitesimal) lightlike convexity of the boundary, that is, the positive semidefiniteness of the $\partial M$ second fundamental form on lightlike directions. More precisely,
@@
The proof relies on a technical result on the equivalence of infinitesimal and local convexity for $\partial M$ in the semi-Finsler case [84, Theor. 1.1], where the anisotropic Chern connection must be used.
It is worth pointing out that, then, $\mathcal{N}$ is a smooth Hausdorff (2n1)(2n-1)-manifold (without boundary), and it can be explicitly described in terms of: (a) the projectivization of the tangent bundle $T\bar{S}$ of any spacelike Cauchy hypersurface $\bar{S}$ in $\bar{M}^{n+1}$ (also used in the case without boundary), (b) the set of inwards lightdirections starting at points in $J^{+}(\bar{S})\cap\partial M$ and (c) the set of outwards lightdirections starting at points in $J^{-}(\bar{S})\cap\partial M$. In particular [84, Corollary 4.14]:
@@
This result opens prospects to extend linking results in Theorem 3.7 to a considerably bigger setting.
@@ 3.3. Singularity theorems
Aazami and Javaloyes [1] proved the first Finslerian singularity theorem by transplanting Penrose’s one. They solved technical issues (as those related with the notion of trapped surfaces or the lack of smootahbility of the exponential map at 0), and clarified the appropiated tools. Indeed, they noticed that the involved elements (lightlike geodesics, focal points, trapped surfaces, Ricci scalar) depended exclusively on the nonlinear connection and cone structure, then using effectively the anisotropic Chern connection. Other classical singularity theorems where extended to Finsler spacetimes focusing on Raychaudhuri equation in [120]. Moreover, weighted Raychaudhuri equations and inequalities were obtained in [115] by using weighted Ricci curvature, then obtaining various singularity theorems in this setting. Such a weighted tensor extends the weighted Ricci curvature introduced for Finsler metrics by Ohta [138].
@@ 4. Lorentz-Finsler applied to classical geometries
Next, the Lorentz-Finsler viewpoint applied to both, Lorentz Geometry, and Finsler one (including the Riemannian case) is developed. The relativistic causal boundary will emerge as a unifying concept for some relevant boundaries in the Riemannian, Finslerian and Lorentzian settings. Most of these applications have already been explained in the review [93]. So, we will make here just a short summary with some updates.
@@ 4.1. Application to causality in Lorentz Geometry
Here, our basic references are [38, 39], apart from the aforementioned [93].
@@ 4.1.1. Stationary spacetimes
which belongs to the Finslerian class of Randers metrics. Notice that this metric is non-reversible (that is, there exist $v\in TS$ such that $F(v)\neq F(-v)$) whenever $\omega\neq 0$, then reflecting an asymmetry between the future and past-pointing directions. As explained in §2.3, $F$ determines a generalized distance $d_{F}$, which is again non-symmetric if $\omega\neq 0$. Then, one can define the forward and backward balls, $B^{+}(x,r)$, $B^{-}(x,r)$ associated to $d_{F}$, whose definition depends on whether they are constructed by using the distance from the center $x$ to the point or the other way round. In terms of these balls, chronological future and pasts can be computed easily and the causality of the spacetime can be characterized in a tidy way, see [38, Theorems 4.3 and 4.4] or [93, §3.4.2]. Summing up:
@@
Any standard stationary spacetime $(M=\mathds{R}\times S,g)$ with associated Fermat metric $F$ (as in (18), (19) above) the spacetime is causally continuous [^17] A causality condition stronger then stable causality (ensured by the existnece of the temporal function $t$), where the sets $I^{\pm}(p)$ vary continuously with $p\in M$. and:
@@ 4.1.2. SSTK spacetimes
(here, $\parallel\cdot\parallel_{0}$ denotes the induced $g_{0}$-norm). In this case, the spacetime is called standard with a spacetransverse Killing vector field or SSTK. For these spacetimes, $t$ is still a temporal function, and the intersection $\{dt=1\}\cap\mathcal{C}$ provides a wind Finsler structure $\Sigma$ on $S$ (Fig. 10). Let us analyze it explicitly (see [39, §3.3] for detailed computations).
@@
However, these data make sense whenever the SSTK condition (20) holds. Indeed, the case $\Lambda=0$ corresponds to the critical case $0\in\Sigma$ (or Zermelo data with $g_{R}(W,W)=1$), so that $F$ turns into a Kropina metric, which is a type of singular Finsler metric (explicit in 32 below). The case $\Lambda<0$ (Zermelo data with $g_{R}(W,W)>1$) corresponds to a strong wind structure. Here, the name wind Riemannian structure can be also used, as the desplaced indicatrix comes from a Riemannian metric.
@@
(see the table in [39, Fig. 2, p. 44] for additional details). As shown exhaustively in [39, Chapter 2], this wind structures admit natural notions on lengths, geodesics, convexity and completeness. In terms of them, the causal properties of the SSTK spacetime can be determined precisely, [39, Chapters 3-5], in particular extending Theorem 4.1, see [39, Th. 5.9].
It is worth pointing out that SSTK structures appear in several relativistic spacetimes, especially, black holes. Indeed, slow Kerr spacetime presents such a Killing $K$, which is ligthlike in the limit stationary hypersurface which bounds the so-called ergosphere (where Penrose process can occur, theoretically allowing for energy to be extracted from the black hole’s rotation). For the specific description of this region and physically related properties, see [98, 50].
@@ 4.2.1. Wind Riemannian manifolds of constant flag curvature
In the previous section we have seen that, in a natural way, Randers metrics with Zermelo data $(g_{R},W)$ turns out Kropina metric if one permits $g_{R}(W,W)=1$ and strong wind Riemannian if $g_{R}(W,W)>1$. Kropina metrics are typically studied as singular Finsler metrics written as a quotient $g_{R}/\omega$ where $\omega$ is a non-vanishing and only the region $\omega>0$ is taken into account. However, the scarcely restrictive condition on strong convexity for the Finslerian indicatrices makes natural that the evolve in a highly non trivial way for relevant classes of Finsler manifolds, in contrast with the Riemannian case (where, in natural coordinates, the indicatrices must be always ellipsoids centerd at 0). Then, it is not surprising that wind Riemannian (or, in general, wind Finslerian) structures may arise as complete solutions for a problem initially posed for Randers (or, in general Finsler) manifolds. In this case, the viewpoint of cone structures provides a natural “non-singular” description of the problem, which may lead to a more comprehensive solution.
Remarkably, this was the case of the problem of classification for Randers metrics of constant flag curvature (CFG). As explained in §2.4.7 flag curvature plays a role with analogies to sectional curvature in (semi-) Riemannian geometry. However, in striking difference with the semi-Riemannian case, the problem of finding the model spaces of constant flag curvature is completely open in Finsler Geometry. Indeed, a milestone in Finslerian Geometry obtained by Bao, Robles and Shen [9] was the classification of such manifolds restricted to the class of Randers metrics. Bao et al. local solution states that the Zermelo data $(g_{R},W)$ of a CFC Randers manifold must be: $g_{R}$ a Riemannian metric of constant curvature $c$ and $W$ any of its Killing vector fields, and, in the case $c=0$, $W$ a homothetic vector field too. For the global result, however, $g(W,W)$ may reach the value $1$ in some of the previous cases, then leading to incomplete solutions. These solutions are inextensible as Randers metrics but can be extended as wind Riemannian ones, as proved in [97]. This leads to the following global result, [97, Theorem 3.12], obtained taking into account the natural notion of completeness for wind Finslerian structures (see [39, Def. 2.45, Prop. 6.6]).
@@
Notice that this result also extends the classification of Kropina metrics by Yoshikawa and Sabau [169].
@@ 4.2.2. Geodesic refocusing
Refocusing is a relevant property in Riemannian Geometry, studied in its own right by Bérard-Bergery in [20], which is related to the problem of manifolds whose geodesics are closed. A Riemannian manifold is called a $Z^{x}$ manifold when all geodesics starting at $x$ return to $x$, and a $Y^{x}_{l}$ manifold if every unit-speed geodesic starting at $x$ arrives at $x$ at time $l>0$. A basic open question is whether there are $Z^{x}$ non $Y^{x}_{l}$ manifolds. Considering two points $x,y$ and geodesics connecting them, analogous definitions $Z^{(x,y)}$, $Y^{(x,y)}_{l}$ follow; they formalize the idea of strong refocusing for the geodesics departing from $x$.
Studying the space of lightlike geodesics $\mathcal{N}$, Kinlaw [104] (see also below Th. 3.7) stressed the importance of refocusing, that is, the property that all the geodesics starting at an event $p$ arrive close to a second one $q$. In the case of a product spacetime $\mathds{R}\times S$, if $S$ is a $Y^{(x,y)}_{l}$ then all the lightlike geodesics starting at $p=(x,0)$ will arrive at $q=(y,l)$, i.e. a strong refocusing holds. Consistently, Chernov, Kinlaw and Sadykov [42] studied Riemannian refocusing.
Recently, Bauermeister [11] has used the Lorentzian viewpoint to partially solve the open question above (among other applications), namely: an analytic Riemannian $Z^{(x,y)}$ is also $Y^{(x,y)}_{l}$ for some $l>0$. This is obtained as a consequence of a stronger Lorentzian result for the class of observer refocusing spacetimes introduced by himself, namely, a globally hyperbolic, observer refocusing spacetime of dimension at least 3, endowed with an analytic metric, is strongly refocusing.
The proof uses the analytic version of the Lorentzian Bernal-Sánchez splitting explained in §3.1.1 (item 7) and other techniques also transplantable to the Lorentz-Finsler setting. Thus, the possibility to obtain new applications of Lorentz-Finsler geometry to Finsler one is open.
@@ 4.3. The causal boundary: a link among the three geometries
The causal boundary of a spacetime, c-boundary for short, is a construction initially proposed by Geroch, Kronheimer and Penrose [69] as an intrinsic alternative to the extrinsic conformal boundary commonly used in General Relativity. In their seminal article, these authors pointed out that their own definition of the topology should be improved. This question became a major issue, that hindered the applicability of the c-boundary. What is more, the c-boundary turned out non-Hausdorff in some simple cases. After the careful study by Flores et al. [62] (see also previous Harris’ [77]), it was clear that a minimal chr-topology should be always present at this boundary and that this topology might be non-Hausdorff. However, it was left open the possibility that, eventually, some authors might prefer to choose a refinement of this topology in certain particular cases.
@@
The links with the Riemannian setting appeared when considering the c-boundary of a product $(\mathds{R}\times S,g=-dt^{2}+g_{0})$. For a Riemannian manifold $(S,g_{0})$, one can consider the following three general boundaries: the basic Cauchy one for incomplete metrics, the celebrated Gromov’s compactification for any complete Riemannian manifold [75, 76] and the Busemann one for Cartan-Hadamard manifolds.
Essentially, Gromov’s identify each $x\in S$ with $d_{x}+\mathds{R}$, that is, the distance function $d_{x}$ to $x$, up to an (additive) constant. Then, $S$ can be regarded as a subset of the space of Lipschitz functions on $S$ up to a constant,. As this space is compact, the compactification of $M$ is attained just by taking the closure therein, being Gromov’s boundary the frontier points.
Busemann compactification provides a sphere at infinite (computable by using Busemann functions) for any Cartan-Hadamard manifold, as proven by Eberlein and O’Neill [48], and turns out equivalent to Gromov’s for them. As also proven by Flores et al., the construction is generalizable to any complete Riemannian manifold, now using Busemann-type functions on diverging curves.
Putting a suitable topology, this generalized Busemann boundary naturally matches with the spacetime c-boundary and may be non-Hausdorff in some patological cases. It is worth pointing out than, in these cases Gromov boundary presents akward properties too. When Busemann’s is Hausdorff, however, Gromov and Busemann boundaries are equivalent and become the natural Rieamannian counterpart to the c-boundary.
@@
Even though the extensions of Gromov and Busemann boundaries to the Finslerian setting can be done, the non-reversible metrics admit a variety of new elements associated with the non-symmetry of the distance. Indeed, one has a forward and backward completion for each one of the three Riemannian completions, which will correspond to the different behavior of the c-boundary of stationary spacetimes to the future and the past.
@@
5. Real world applications
Finsler and Lorentz-Finsler Geometry have been applied to study moving objects and wave propagation in many situations. These include Zermelo’s navigation problem (the motion of a ship in a current or a zeppelin in the air), [9, 39, 99, 119], general anisotropic waves [92], or more specific ones, such as sound waves [71, 72], wildfire fronts [94, 118] and seismic waves [4, 168], the latter including refraction between different layers of the Earth [32, 155].
The specific Lorentz-Finsler viewpoint was stressed in the first version of [39]. There it is shown the possibility to model, in a unified and non-singular way, the motion of objects whose velocities yield a wind-Finsler structure by using a cone structure (as in Fig. 7). Then, the systematic application of cone structures and Lorentz-Finsler metrics in [92], and subsequent applications as [94], showed a number of advantages, which will appear next:

### v2

@@ 3.2.2. Causal simplicity and the case with timelike boundary
In the case of spacetimes with timelike boundary seen in §3.1, the recent reference [91] considers both, Lorentz-Finsler metrics and spacetimes with timelike boundary (although the outcomes are equal in the Lorentz and the Lorentz-Finsler cases, the proof in the latter becomes subtler). Indeed, both causal simplicity and Hausdorffness of $\mathcal{N}$, are ensured by the (infinitesimal) lightlike convexity of the boundary, that is, the positive semidefiniteness of the $\partial M$ second fundamental form on lightlike directions. More precisely,
@@
The proof relies on a technical result on the equivalence of infinitesimal and local convexity for $\partial M$ in the semi-Finsler case [91, Theor. 1.1], proved by using naturally the anisotropic Chern connection.
It is worth pointing out that, then, $\mathcal{N}$ is a smooth Hausdorff (2n1)(2n-1)-manifold (without boundary), and it can be explicitly described in terms of: (a) the projectivization of the tangent bundle $T\bar{S}$ of any spacelike Cauchy hypersurface $\bar{S}$ in $\bar{M}^{n+1}$ (also used in the case without boundary), (b) the set of inwards lightdirections starting at points in $J^{+}(\bar{S})\cap\partial M$ and (c) the set of outwards lightdirections starting at points in $J^{-}(\bar{S})\cap\partial M$. In particular [91, Corollary 4.14]:
@@
This result opens prospects to extend some of the aforementioned results to a considerably bigger setting.
@@ 3.3. Singularity theorems
Aazami and Javaloyes [1] proved the first Finslerian singularity theorem by transplanting Penrose’s one. They solved technical issues (as those related with the notion of trapped surfaces or the lack of smoothability of the exponential map at 0), and clarified the appropiated tools. Indeed, they noticed that the involved elements (lightlike geodesics, focal points, trapped surfaces, Ricci scalar) depended exclusively on the nonlinear connection and cone structure, then using effectively the anisotropic Chern connection. Other classical singularity theorems where extended to Finsler spacetimes focusing on Raychaudhuri equation in [130]. Moreover, weighted Raychaudhuri equations and inequalities were obtained in [124] by using weighted Ricci curvature, then obtaining various singularity theorems in this setting. Such a weighted tensor extends the weighted Ricci curvature introduced for Finsler metrics by Ohta [147].
@@ 4. Lorentz-Finsler applied to classical geometries
Next, the Lorentz-Finsler viewpoint applied to both, Lorentz Geometry, and Finsler one (including the Riemannian case) is developed. In particular, the relativistic causal boundary will emerge as a unifying concept for some relevant boundaries in the Riemannian, Finslerian and Lorentzian settings. Most of these applications have already been explained in the review [102]. So, we will make here just a short summary with some updates.
@@ 4.1. Application to causality in Lorentz Geometry
Consistently with the Hamilton-Jacobi viewpoint explained in §2.5.4, next the causal futures and pasts in a stationary spacetime can be computed by using a convenient cone triple $(\Omega=dt,T,F)$, where $F$ is a Finsler metric of Randers metric which is used as a cost function. Then, the causal futures and past of a subset are described in terms of a Randers distance function, whose possible non-symmetry reflects the asymmetry of the time $t$ (the map $t\mapsto-t$ is not an isometry). This leads to a precise characterization of the causality of these spacetimes, including the position in the causal ladder. Moreover, the applications will be extended to the class of $SSTK$ spacetimes by dropping the hypothesis that $T$ in the cone triple is timelike. Here, our basic references are [40, 41], apart from the aforementioned [102].
@@ 4.1.1. Stationary spacetimes
which belongs to the Finslerian class of Randers metrics. Notice that this metric is non-reversible (that is, there exist $v\in TS$ such that $F(v)\neq F(-v)$) whenever $\omega\neq 0$, then reflecting an asymmetry between the future and past-pointing directions. As explained in §2.3, $F$ determines a generalized distance $d_{F}$, which is again non-symmetric if $\omega\neq 0$. Then, one can define the forward and backward balls, $B^{+}(x,r)$, $B^{-}(x,r)$ associated to $d_{F}$, whose definition depends on whether they are constructed by using the distance from the center $x$ to the point or the other way round, respectively. In terms of these balls, chronological future and pasts can be computed easily and the causality of the spacetime can be characterized in a tidy way, see [40, Theorems 4.3 and 4.4] or [102, §3.4.2]. Summing up:
@@
Any standard stationary spacetime $(M=\mathds{R}\times S,g)$ with associated Fermat metric $F$ (as in (21), (22)) is causally continuous [^16] A causality condition stronger then stable causality (ensured by the existence of the temporal function $t$), where the sets $I^{\pm}(p)$ vary continuously with $p\in M$, [137]. and:
@@ 4.1.2. SSTK spacetimes
(here, $\parallel\cdot\parallel_{0}$ denotes the induced $g_{0}$-norm). In this case, the spacetime is called standard with a spacetransverse Killing vector field or SSTK. For these spacetimes, $t$ is still a temporal function, and the intersection $\{dt=1\}\cap\mathcal{C}$ provides a wind Finsler structure $\Sigma$ on $S$ (Fig. 10). Let us analyze it explicitly (see [41, §3.3] for detailed computations).
@@
However, these data make sense whenever the SSTK condition (23) holds. Indeed, the case $\Lambda=0$ corresponds to the critical case $0\in\Sigma$ (or Zermelo data with $g_{R}(W,W)=1$), so that $F$ turns into a Kropina metric, which is a type of singular Finsler metric (explicit in 35 below). The case $\Lambda<0$ (Zermelo data with $g_{R}(W,W)>1$) corresponds to a strong wind structure. Here, the name wind Riemannian structure can be also used, as the desplaced indicatrix comes from a Riemannian metric.
@@
(see the table in [41, Fig. 2, p. 44] for additional details). As shown exhaustively in [41, Chapter 2], this wind structures admit natural notions on lengths, geodesics, convexity and completeness. In terms of them, the causal properties of the SSTK spacetime can be determined precisely, [41, Chapters 3-5], in particular extending Theorem 4.1, see [41, Th. 5.9].
It is worth pointing out that SSTK structures appear in several relativistic spacetimes, especially, black holes. Indeed, slow Kerr spacetime presents such a Killing $K$, which is ligthlike in the limit stationary hypersurface which bounds the so-called ergosphere [^19] The ergosphere has relevant physical properties. Indeed, Penrose process can occur, theoretically allowing for energy to be extracted from the black hole’s rotation.. For the specific description of this region and physically related properties, see [106, 54].
@@ 4.2.1. Wind Riemannian manifolds of constant flag curvature
In the previous section we have seen that, in a natural way, Randers metrics with Zermelo data $(g_{R},W)$ turns out Kropina metrics if one permits $g_{R}(W,W)=1$ and strong wind Riemannian if $g_{R}(W,W)>1$. Kropina metrics are typically studied as singular Finsler metrics written as a quotient $g_{R}/\omega$, where $\omega$ is a non-vanishing 1-form and only the region $\omega>0$ is taken into account. In Finslerian Geometry becomes natural the evolution of the centroid of the indicatrix, in contrast with the Riemannian case where, in natural coordinates, the indicatrices must be always ellipsoids centered at 0. Then, it is not surprising that wind Riemannian (or, in general, wind Finslerian) structures may arise as complete solutions for a problem initially posed for Randers (or, in general Finsler) manifolds. In this case, the viewpoint of cone structures provides a natural “non-singular” description of the problem, which may lead to a more comprehensive solution.
Remarkably, this was the case of the problem of classification for Randers metrics of constant flag curvature (CFG). As explained in §2.4.7 flag curvature plays a role with analogies to sectional curvature in (semi-) Riemannian geometry. However, in striking difference with the semi-Riemannian case, the problem of finding the model spaces of constant flag curvature is completely open in Finsler Geometry. Indeed, a milestone in Finslerian Geometry obtained by Bao, Robles and Shen [11] was the classification of such manifolds restricted to the class of Randers metrics. Bao et al. local solution states that the Zermelo data $(g_{R},W)$ of a CFC Randers manifold must be: $g_{R}$ a Riemannian metric of constant curvature $c$ and $W$ any of its Killing vector fields, and, in the case $c=0$, $W$ a homothetic vector field too. For the global result, however, $g(W,W)$ may reach the value $1$ in some of the previous cases, then leading to incomplete solutions. These solutions are inextensible as Randers metrics but can be extended as wind Riemannian ones, as proved in [105]. This leads to the following global result, [105, Theorem 3.12], obtained taking into account a natural notion of completeness for wind Finslerian structures (see [41, Def. 2.45, Prop. 6.6]).
@@
Such a result also extends the classification of Kropina metrics by Yoshikawa and Sabau [186].
@@ 4.2.2. Geodesic refocusing
Refocusing is a relevant property in Riemannian Geometry, studied in its own right by Bérard-Bergery in [22], which is related to the problem of manifolds whose geodesics are closed. A Riemannian manifold is called a $Z^{x}$ manifold when all geodesics starting at $x$ return to $x$, and a $Y^{x}_{l}$ manifold if every unit-speed geodesic starting at $x$ arrives at $x$ at time $l>0$. A basic open question is whether there are $Z^{x}$ non $Y^{x}_{l}$ manifolds. Considering two points $x,y$ and geodesics connecting them, analogous definitions $Z^{(x,y)}$, $Y^{(x,y)}_{l}$ follow and they formalize the idea of strong refocusing for the geodesics departing from $x$.
Studying the space of lightlike geodesics $\mathcal{N}$, Kinlaw [112] (see also below Th. 3.7) stressed the importance of refocusing, that is, the property that all the geodesics starting at an event $p$ arrive close to a second one $q$. In the case of a product spacetime $\mathds{R}\times S$, if $S$ is a $Y^{(x,y)}_{l}$ then all the lightlike geodesics starting at $p=(x,0)$ will arrive at $q=(y,l)$, i.e. a strong refocusing holds. Consistently, Chernov, Kinlaw and Sadykov [45] studied Riemannian refocusing. Recently, Bauermeister [13] has used the Lorentzian viewpoint to study the open question on $Z^{(x,y)}$, $Y^{(x,y)}_{l}$ above in the analytic case. The proofs use techniques transplantable to the Lorentz-Finsler setting (as the splitting explained in §3.1.1, item 7), opening the possibility to obtain new applications of Lorentz-Finsler geometry.
@@ 4.3. The causal boundary: a link among the three geometries
The causal boundary of a spacetime, c-boundary for short, is a construction initially proposed by Geroch, Kronheimer and Penrose [73] as an intrinsic alternative to the extrinsic conformal boundary commonly used in General Relativity. In their seminal article, these authors pointed out that their own definition of the topology should be improved. This question became a major issue, that hindered the applicability of the c-boundary. What is more, the c-boundary turned out non-Hausdorff in some simple cases, making it awkward. After the careful study by Flores et al. [68] (see also previous Harris’ [82]), it was clear that a minimal chr-topology should be always present at this boundary and that this topology might be non-Hausdorff. Anyway, it was left open the possibility that, eventually, some authors might prefer to choose a refinement of this topology in certain particular cases.
@@
Following [68], the links with the Riemannian setting appeared when considering the c-boundary of a product $(\mathds{R}\times S,g=-dt^{2}+g_{0})$. For a Riemannian manifold $(S,g_{0})$, one can consider the following three general boundaries: the basic Cauchy one for incomplete metrics, the celebrated Gromov’s compactification for any complete Riemannian manifold [80, 81] and the Busemann one for Cartan-Hadamard manifolds.
Essentially, Gromov’s identify each $x\in S$ with $d_{x}+\mathds{R}$, that is, the distance function $d_{x}$ to $x$, up to an (additive) constant. Then, $S$ can be regarded as a subset of the space of Lipschitz functions on $S$ up to a constant. As this space is compact, the compactification of $M$ is attained just by taking the closure therein, being Gromov’s boundary the frontier points.
Busemann compactification provides a sphere at infinite (computable by using Busemann functions) for any Cartan-Hadamard manifold, as proven by Eberlein and O’Neill [55]. It turns out equivalent to Gromov’s for these spaces. As proven in [68], the construction is generalizable to any complete Riemannian manifold, now using Busemann-type functions on diverging curves. Using a suitable topology, this generalized Busemann boundary naturally matches with the spacetime c-boundary and may be non-Hausdorff in some patological cases. It is worth pointing out than, in these cases Gromov boundary presents akward properties too. When Busemann’s is Hausdorff, however, Gromov and Busemann boundaries are equivalent and become the natural Rieamannian counterpart to the c-boundary.
@@
The extensions of Gromov and Busemann boundaries to the reversible Finslerian setting have a similar fashion. However, non-reversible Finsler metrics admit a variety of new elements associated with the non-symmetry of the distance. Indeed, one has a forward and backward completion for each one of the three Riemannian completions, which will correspond to the different behavior of the c-boundary of stationary spacetimes to the future and the past.
@@
5. Applications in Everyday Physics
Finsler and Lorentz-Finsler Geometry have been applied to study moving objects and wave propagation in many situations. These include Zermelo’s navigation problem (the motion of a ship in a current or a zeppelin in the air), [11, 41, 107, 128], general anisotropic waves [101], or other specific ones, such as sound waves [75, 76], wildfire fronts [103, 127] and seismic waves [5, 185], the latter including refraction between different layers of the Earth [34, 170].
The specific Lorentz-Finsler viewpoint was stressed since the first version of [41]. There it is shown the possibility to model, in a unified and non-singular way, the motion of objects whose velocities yield a wind-Finsler structure by using a cone structure (as in Fig. 7). Then, the systematic application of cone structures and Lorentz-Finsler metrics in [101], and subsequent applications as [103], showed a number of advantages, which will appear next: